# SLOctolyzer: Fully automatic analysis toolkit for segmentation and feature extracting in scanning laser ophthalmoscopy images

### v1

@@ 3.3 Execution time
Table 6 shows the execution time of SLOctolyzer’s segmentation module and entire pipeline. Because the segmentation module resizes all images to a common resolution of 768 $\times$ 768, the segmentation execution time across all image types and resolutions are approximately equal. For binary vessel segmentation, it takes approximately 4.5 s/image, for the lightweight fovea detection model it takes approximately 0.6 s/image, and for artery-vein-optic disc detection it takes approximately 3.0 s/image. Thus, SLOctolyzer’s segmentation module using a standard laptop CPU takes approximately 11.1 $\pm$ 0.5 s/image. Measuring the execution time of the entire pipeline on a standard laptop CPU took 18.9 $\pm$ 1.0s for a macula-centred SLO image at 768 $\times$ 768 resolution, 28.6 $\pm$ 1.0s for a disc-centred SLO image at 768 $\times$ 768 resolution and 111.0 $\pm$ 3.5s for a disc-centred SLO image at 1536 $\times$ 1536 resolution. Disc-centred SLO images are more time intensive for feature measurement as measurements are made across the whole image, and for zones B and C across the vessel, artery and vein segmentation maps separately.
@@
Table 6: Average (standard deviation) execution time in seconds of the segmentation module and SLOctolyzer’s entire pipeline, for SLO images of different location and size.
@@ 4 Discussion
We have developed a fully automatic method, SLOctolyzer, for processing macula-centred and disc-centred SLO images which typically accompany OCT data but are often discarded in research. SLOctolyzer converts a raw SLO image into clinically relevant retinal vascular parameters, including fractal dimension, tortuosity, and vessel calibre.
@@
SLOctolyzer performed well on unseen data, but we observed a performance drop against an external test set featuring severe retinal pathology. SLOctolyzer was designed for modelling the retinal vasculature during systemic health change, so there was understandably a noticeable decrease in performance as the subjective extent of pathology increased. However, we believe SLOctolyzer’s additional functionality to correct any annotations and re-compute measurements will prove very useful for cases of severe retinal pathology. Regardless, we envision SLOctolyzer being a very useful tool for SLO image analysis, particularly as the nascent field of oculomics increases rapidly 41, since the populations being studied often have grossly normal retinae.
@@
SLOctolyzer’s segmentation module had good reproducibility, with the residuals across all metrics for vessels, arteries and veins exhibiting no pattern and being centred around 0. In particular, the residuals for measurements of global and local calibre (in microns) appear to show proportionately small error in comparison to what is generally considered as clinically significant. Meta-analyses have found vessel calibre difference of 10–15 microns between mid- and old-age, with 20 micron increase in venule calibre associated with increased risk of stroke by 1.15 42.
@@
The distributions of measurement error at the eye-level for fractal dimension, vessel density and local/global calibre were found to be reasonable given the homogeneous nature of the i-Test cohort studied for reproducibility — with tortuosity density and large-vessel calibre having higher variability. We observed that small differences in the segmentation maps can have an inordinate impact on tortuosity density (supplementary Fig. S10(A)). This highlights the apparent sensitivity of tortuosity, regardless of whether the segmentation looks qualitatively different. While segmentation error contributes to this error, this observation adds to a potential problem surrounding standardisation of retinal vascular parameters 43. For large-vessel calibre, the largest outliers for large-vessel calibre came from the same eye, which showed a major arteriovenous crossings at the edge of the image causing a major artery-vein miss-classification. Moreover, large-vessel calibre metrics like AVR/CRAE/CRVE are typically reserved for disc-centred SLO images as these show the major vessels in full. Thus, the unregistered nature of the macula-centred SLO pairs, with major vessels appearing toward the edge of the image, also contributed to the measurement error observed (supplementary Fig. S10(B)). We hope in the future we may also assess SLOctolyzer’s reproducibility for paired, disc-centred SLO images.
On a standard laptop CPU, SLOctolyzer can segment an SLO image in around 11 seconds. For a standard 768 $\times$ 768 resolution image, the entire pipeline takes 19 seconds for macula-centred images and 29 seconds for disc-centred images. It can be used interactively or for batch processing (supplementary Fig. S8), converting raw SLO images into reproducible vascular metrics and outputting intermediary segmentation masks for quality inspection (supplementary Fig. S9).
@@
There were some limitations associated with this work. Firstly, the SLO images used to build SLOctolyzer’s segmentation module were sourced from only one manufacturer, Heidelberg Engineering. Thus, we do not know the extent of performance from SLO images from other devices. Furthermore, most of the images we used were from individuals recruited to studies of systemic health and did not have obvious retinal pathology. We therefore expect the segmentation module to not be entirely robust to cases of severe retinal pathology, as seen by the performance against the RAVIR dataset (Table 4), which featured severe retinal pathology. However, SLOctolyzer’s additional functionality allows for segmentation error to be corrected manually.
@@
A technical limitation of the work is that SLOctolyzer’s segmentation module is made up of three distinct models. The rationale for this was clear at the time of model construction: we could leverage an initial binary vessel segmentation model to help us manually annotate the arteries and veins for building an artery-vein detection model, a much more challenging task, with the fovea detection model utilising all available images due to the speed of foveal pit manual annotation. However, these models now provide a quick and consistent way to annotate large datasets which, after inspection and correction by expert graders, could be used to build a much more robust segmentation model.
Finally, our datasets were of macula-centred and disc-centred SLO images, and thus we do not know its performance at different locations on the retina.
@@
We intend to improve SLOctolyzer’s segmentation module in future work by constructing a single model to perform all segmentation tasks and refine the dataset on which it will be trained and evaluated on. This will ideally include more diverse SLO images, such as those related to retinal pathology, capturing different regions of the retina, and from multiple imaging devices and manufacturers. Improving SLOctolyzer’s ability to generalise to other SLO imaging modalities, such as Optos ultra-wide field SLO 44, or Heidelberg’s more recent multicolour SLO 45, would further increase the scope of this open-source pipeline.
@@ 5 Conclusion
We have developed a fully automatic method for extracting useful information from SLO images, including fractal dimension, tortuosity, and vessel calibre. SLOctolyzer, the first open-source analysis toolkit for macula- and disc-centered SLO images, addresses the lack of open-source tools for this image modality. It leverages existing OCT imagesets to provide clinically meaningful retinal vascular parameters from the SLO images which are captured in parallel to OCT acquisition, but are often overlooked. Available on GitHub, SLOctolyzer requires no specialist training or proprietary software and supports batch processing and manual correction. We hope this tool will encourage the research community to reconsider the value of SLO images, facilitate collaboration and help to standardise retinal vascular parameters in SLO images and retinal imaging more generally.
@@ 6 Conflicts of Interest
Supplementary Fig. S2 shows some exemplary SLO images selected from the i-Test and FutureMS cohorts for building the artery-vein-optic disc segmentation model. The total set of 30 SLO images were selected to provide a variety of image-features during training and evaluation, such as blur, illumination, and contrast artefacts, as well as abnormal retinal features such as vessel tortuosity or optic nerve head atrophy.
@@
Figure S2: Selection of the data used to build the artery-vein-optic disc segmentation model, highlight features of interest in the image which are advantageous for diverse training, such as blur (A), contrast artefacts (B,E,H), illuminations (D, F, G), vessel tortuosity (C) and optic nerve head atrophy (H).
Repeatability of fovea manual detection
@@
As part of the manual detection of the fovea coordinate on SLO images, author J.B. manually selected the fovea coordinate for all cohorts but the i-Test one (as we could cross-reference the fovea on the SLO using the fovea-centred OCT B-scan). As a measure of rater variability, author J.B. re-selected the fovea coordinate for all 516 SLO images two months after the initial manual detection, and compared the $(x,y)$-positioning using mean absolute error and intra-class correlation. Like the training pipeline for the fovea detection model, fovea segmentation masks were generated for each batch of coordinates and compared using the area under the receiver operating characteristic curve (AUC) and dice coefficient. Evaluation was performed after all images were resized to a common pixel resolution of 768 $\times$ 768 for interpretable comparison.
@@
The repeatability results are listed below in supplementary Table S1. Results for the i-Test cohort represent inter-observer variability (manual detection against accurate ground truth data), and results for the remaining four cohorts of data represent intra-observer variability. Intra-observer variability has an excellent average intra-class correlation of 0.9934 and a strong intra-class correlation for inter-observer variability of 0.8206.
@@
Others 0.9640 0.9287 0.9998 0.9870 3.3921 3.0000
i-Test 0.9441 0.8892 0.7720 0.8692 4.7527 4.5161
Table S1: Repeatability of detecting the fovea on the i-Test dataset (which act as ground truth) and the remaining cohorts (which measure variability in the rater (J.B.). AUC, area under the receiver operating characteristic curve; ICC, intra-class correlation for single fixed raters; MAE, mean absolute error (in pixels)
@@
    Alternating rule: It’s common to see arteries and veins appear on the fundus in an alternative nature, for efficient oxygen transport 46.
@@
    Thickness rule: Veins are typically thicker than arteries 40.
@@
    Brightness rule: Veins are typically darker than arteries 47.
@@
Figure S3: An exemplary SLO image with annotations to describe each of the four core rules for artery-vein classification.
A protocol was created to ensure consistent and comparable segmentations by each grader using ITK-Snap 22. After application of the binary vessel segmentation model to detect all the vessels in each SLO image, the core segmentation rules the graders followed are listed below:
@@
    Remove false positive pixels, i.e. retinal tissue which have been classed as vessel.
@@
Figure S4: An exemplary SLO image (A) with a raw binary vessel mask (B) superimposed in green with annotations to describe each segmentation rules followed during annotation.
@@
Grader 1 vs. Grader 2 0.9898 0.9597 0.9703 0.9442
Grader 1 vs. Grader 3 0.9906 0.9526 0.9602 0.9480
Grader 2 vs. Grader 3 0.9887 0.9468 0.9596 0.9379
@@
Supplementary Table S3 show the qualitative grading results from the clinical ophthalmologist (I.M.) for the manual annotations. All SLO images were graded as having ‘good quality’ or above, except for one, and there was no annotation which received a bad or worse rating. Across the image quality categories, a total of 19 SLO images were marked as ‘very good’, 10 as ‘good’ and 1 as ’okay’.
@@
One of our image graders (author J.B.) identified any major errors in the original segmentation labels and corrected them using ITK-Snap 22. Supplementary Fig. S6 shows an SLO image with the most significant error (A), the SLO image with the original labelling overlaid (B), and the SLO image with the corrected version overlaid (C).
@@
Figure S6: SLO image (A), original manual annotation superimposed 10 (B) and corrected manual annotation (C). Here, many of the vessels in the macula were all assumed to be venules in the original annotation.
Image Quality Original 10 Corrected Both Very good ($N=6$) 0 3 3 Good ($N=4$) 0 2 2 Okay ($N=10$) 0 7 3 Bad ($N=3$) 0 2 1 Total ($N=23$) 0 14 9 Label Artery-Vein classification Original 10 VG: 6, G: 8, O: 6, B: 2, VB: 1 Corrected VG: 10, G: 11, O: 2, B: 0, VB: 0
Table S4: Qualitative adjudication between the original and corrected manual annotations for artery and vein classification in the RAVIR dataset 10, from a clinical ophthalmologist (I.M.). VG, very good; G, good; O, okay; B, bad; VB, very bad.
@@
Supplementary Fig. S7 shows the key differences in global and local vessel calibre. Global vessel calibre is the ratio between vessel pixels and skeletonised vessel pixels (number of yellow pixels divided by blue pixels) (B). This provides a coarse measure of calibre. Local vessel calibre measures the diameter of each vessel segment by traversing it’s length and measuring the diameter of the largest circle which fits entirely within it. A vessel segments’ calibre is then the average value of the list of circle diameters collected as the vessel segment is traversed. Local vessel calibre is thus the overall average of all vessel segment calibres, hence a more granular approach to assess vessel calibre.
Figure S7: SLO binary vessel map (A), binary vessel map with skeleton overlaid in dark blue (B) and binary vessel map with individual vessel segments overlaid in an array of colours (C).
@@
Supplementary Fig. S8 shows the core steps to run the SLOctolyzer pipeline. SLOctolyzer can be run either from an integrated development environment (IDE), such as Jupyter, Spyder or Visual Studio Code, or directly from the terminal. All that is required is the path to the directory of images to be analysed, and where the results are to be stored in a simple configuration text file. A process log for the user is outputted as the pipeline processes each image in turn, utilising GPU acceleration if available. An optional spreadsheet with additional metadata such as the conversion factor from pixel space to physical space can be provided (in microns-per-pixel), so that measurements in pixel units can be converted into microns.
Figure S8: Summary of the process of running SLOctolyzer on a batch ($n=3$) of images. Directory of input SLO images to be analysed, with an optional spreadsheet which allows manual data entry of pixel length scale, eye type and location (A). Spreadsheet with optional metadata to accompany SLO images during pipeline. Note that missing information is supported (B). Configuration file specifying the file paths to images and where results will be saved (C). Demonstration of running SLOctolyzer via JupyterLab or via the terminal (D).
@@
Figure S9: Summary of the output of running SLOctolyzer on a batch ($n=3$) of images. Directory of output folders for each of the three image files analysed (A). Output files for the ‘opticdisc_768’ exemplar SLO image, including the original SLO image, segmentation masks, composite image with segmentations superimposed and a results file with metadata and feature measurements (B). A directory with composite images of SLO and segmentation masks per image file for quick inspection of segmentation quality (C). A composite results file with the metadata and features of each image file processed, stored row-wise to facilitate downstream statistical analysis (D).
@@
Supplementary Fig. S10 shows two examples which had the second largest outliers for tortuosity density and large-vessel calibre (AVR/CRAE/CRVE). In supplementary Fig. S10(A), while there is no qualitative difference in the binary vessel segmentation, tortuosity density is quite large. This is likely due to vessel disconnectedness between the segmentations (purple arrows). Nevertheless, the local vessel calibre and fractal dimension residuals are very low.

### v2

@@ 3.3 Execution time
Table 6 shows the execution time of SLOctolyzer’s segmentation module and pipeline. Because the segmentation module resizes all images to a common image resolution of 768 $\times$ 768 pixels, the segmentation execution time across different locations and image resolutions are approximately equal. For binary vessel segmentation, it takes approximately 4.5 s/image, for the fovea detection model it takes approximately 0.6 s/image, and for AVOD detection it takes approximately 3.0 s/image. Thus, SLOctolyzer’s segmentation module using the laptop CPU takes approximately 11.1 $\pm$ 0.5 s/image. Measuring the execution time of the entire pipeline on the laptop CPU took 18.9 $\pm$ 1.0s for a macula-centred SLO with image resolution 768 $\times$ 768 pixels, 28.6 $\pm$ 1.0s for a disc-centred SLO with image resolution 768 $\times$ 768 pixels and 111.0 $\pm$ 3.5s for a disc-centred SLO with image resolution 1536 $\times$ 1536 pixels. Disc-centred SLO images are more time intensive as measurements are made across the whole image, and for zones B and C across the all-vessel, artery and vein segmentation maps separately.
@@
Table 6: Average (standard deviation) execution time in seconds of the segmentation module and SLOctolyzer’s entire pipeline for SLO images of different location and image resolution.
@@ 4 Discussion
We have developed a fully automatic method, SLOctolyzer, for processing macula-centred and disc-centred SLO images which typically accompany OCT data but are often discarded in research. SLOctolyzer converts a raw SLO image into reproducible and clinically meaningful retinal vascular parameters, including fractal dimension, tortuosity, and vessel calibre.
@@
SLOctolyzer performed well on unseen data, but we observed a performance drop against an external test set featuring severe retinal pathology. SLOctolyzer’s segmentation module was trained on datasets primarily related to systemic health and was thus designed for modelling the retinal vasculature in these disease contexts, so there was understandably a noticeable decrease in performance as the subjective extent of pathology increased. However, we believe SLOctolyzer’s additional functionality to correct any erroneous segmentation maps and re-compute measurements will prove very useful for cases of severe retinal pathology. Regardless, we envision SLOctolyzer being a very useful tool for SLO image analysis in systemic disease, particularly as the nascent field of oculomics increases rapidly 42 and the populations being studied often have grossly normal retinae.
@@
SLOctolyzer’s segmentation module had good reproducibility, with the residuals across all metrics for vessels, arteries and veins exhibiting no pattern and being centred around 0. In particular, the residuals for measurements of global and local calibre (in microns) appear to show proportionately small error in comparison to what is generally considered as clinically significant. Meta-analyses have found vessel calibre difference of 10 – 15 microns between mid- and old-age, with a 20 micron increase in venule calibre associated with increased risk of stroke by 1.15 43.
@@
The distributions of measurement error at the eye-level for fractal dimension, vessel density and local/global calibre were found to be reasonable given the homogeneous nature of the i-Test cohort studied for reproducibility — with tortuosity density and large-vessel calibre having higher variability. We observed that small differences in the segmentation maps can have an inordinate impact on tortuosity density (supplementary Fig. S13(A)). This highlights the apparent sensitivity of tortuosity density, regardless of whether the segmentation looks qualitatively different. Segmentation error does contribute to this error, as arteriovenous crossings result in artery-vein map disconnectedness which can have an inordinate impact on generating individual vessel segments to measure tortuosity on. This observation adds to a potential problem surrounding standardisation of retinal vascular parameters 44.
For large-vessel calibre, the largest outliers came from the same eye, which showed a major arteriovenous crossing at the edge of the image causing an artery-vein miss-classification. Moreover, large-vessel calibre metrics like AVR/CRAE/CRVE are typically reserved for disc-centred SLO images as these show the major vessels in full as they branch from the optic nerve head. Thus, the unregistered nature of the macula-centred SLO pairs, combined with major vessels potentially cropped toward the edge of the image, likely contributed to the measurement error (supplementary Fig. S13(B)). We hope in the future we may also assess SLOctolyzer’s reproducibility for paired, disc-centred SLO images.
[…]
There were some limitations associated with this work. Firstly, the SLO images used to build SLOctolyzer’s segmentation module were sourced from only one manufacturer (Heidelberg Engineering) and we do not know SLOctolyzer’s performance on SLO images from other devices. Moreover, our datasets only contained macula- and disc-centred SLO images captured at a field of view of 30-degrees, and we did not explicitly test the pipeline’s performance at different locations or fields of view in the posterior pole.
@@
Furthermore, most of the images were from individuals recruited to studies on systemic disease which did not have obvious retinal pathology and we therefore expect the segmentation module to not be entirely robust to these cases. Finally, these individuals were of white ethnicity, and so we do not know the pipeline’s performance across varying retinal pigmentation. However, SLOctolyzer’s additional functionality allows for any unexpected segmentation errors to be corrected manually via ITK-Snap 23.
@@
A technical limitation of the work is that SLOctolyzer’s segmentation module is made up of three distinct models. The rationale for this was clear at the time of model construction: we could leverage an initial binary vessel segmentation model to help us manually annotate the arteries and veins for building an artery-vein detection model, a much more challenging task, with the fovea detection model utilising all available images due to the speed of ground truth labelling. However, these models now provide a quick and consistent way to annotate large datasets which, after inspection and correction by expert graders, could be used to build a much more robust segmentation model.
@@
Thus, we intend to improve SLOctolyzer’s segmentation module in future work by constructing a single model to perform all segmentation tasks, and include a larger and more diverse dataset on which it will be trained and evaluated. These will include examples related to retinal pathology, capturing different regions and fields of view of the retina, and from multiple imaging devices and manufacturers.
We anticipate SLOctolyzer’s relevancy for processing 820nm SLO images which traditionally accompany OCT image capture using Heidelberg Engineering imaging devices. However, further exploration will be conducted on its generalisability to other SLO imaging modalities, such as Optos ultra-wide field SLO 45, or Heidelberg’s more recent multicolour SLO 46.
@@ 5 Conclusion
We have developed a fully automatic method for extracting useful information on the retinal vessels from SLO images, including fractal dimension, tortuosity, and vessel calibre. SLOctolyzer, the first open-source analysis toolkit for macula- and disc-centred SLO images, addresses the lack of open-source tools for this image modality. It can leverage existing OCT datasets to provide reproducible and clinically meaningful retinal vascular parameters from the SLO images which are captured in parallel to OCT acquisition, but are often overlooked. Available on GitHub, SLOctolyzer requires no specialist training or proprietary software and supports batch processing and manual correction. We hope this tool will encourage the research community to reconsider the value of SLO images, facilitate collaboration and help to standardise retinal vascular parameters in SLO images.
@@ Acknowledgements
Data collection for the Retinal Imaging sub study in PREVENT Dementia was supported in part by the Alzheimer’s Drug Discovery Foundation (project no. GDAPB-201808-2016196, “Delivering novel neuro-retinal biomarkers for the early diagnosis of Alzheimer’s disease”); NHS Lothian R&D; British Heart Foundation Centre for Research Excellence Award III (RE/18/5/34216); a University of Edinburgh Innovation Initiative Grant award; a ARUK Scotland Network Centre Equipment award; the Edinburgh & Lothians Health Foundation. Data collection for i-Test was funded by the Wellcome Leap In Utero programme. These funding sources were not involved in designing, conducting, or submitting this work.
@@ 6 Conflicts of Interest
Supplementary Fig. S2 shows some exemplary SLO images selected from the i-Test and FutureMS cohorts for building the artery-vein-optic disc segmentation model. The total set of 30 SLO images were selected to provide a variety of image-features during training and evaluation, such as blur, non-uniform illumination, and contrast artefacts, as well as abnormal retinal features such as vessel tortuosity or optic nerve head atrophy.
@@
Figure S2: Selection of the data used to build the artery-vein-optic disc segmentation model, highlighting features of interest in the image which are advantageous for diverse training, such as blur (A), contrast artefacts (B,E,H), non-uniform illumination (D, F, G), vessel tortuosity (C) and optic nerve head atrophy (H).
Manual detection of the fovea
Examples of detecting the fovea
As part of the manual detection of the fovea coordinate on SLO images, author J.B. manually selected the fovea coordinate for all cohorts but the i-Test one.
For en face SLO fovea detection the pixel coordinate representing the fovea pit was detected as the centre of a small, dark and circular region representing the depression on the retinal tissue (sometimes showing a hyperreflective spot at the centre of this region, representing the foveal pit). This is typically seen in the centre of the macula-centred SLO image, and at the centre-row but skewed to the far right- or left-column for disc-centred images, depending on the eye’s laterality. To aid detection, the small arterioles and venules branching from the major vessels were used to guide the centre of the macula and thus the fovea pit for more challenging cases. Supplementary Fig. S3 shows some examples of en face SLO images with their foveal pit identified by red arrows, with supporting text to describe the detection in each case.
Figure S3: Examples of foveal pit detection for en face SLO images, with text overlaid to describe the detection process in each case.
For a fovea-centred OCT B-scan, we used Choroidalyzer 16 to define the foveal pit as a single pixel coordinate is at the point of deepest depression in the B-scan. This appears as a dip in the retina with a small hyperreflective region in the centre of this depression, often aligned with a ridge formed at the photoreceptor layer. Fortunately, this detection process is deterministic through Choroidalyzer, and was cross-referenced to the corresponding en face SLO image as we know the transversal position of the fovea on the B-scan and the exact location of acquisition of the B-scan on the SLO image. Supplementary Fig. S4 shows some examples of cross-sectional OCT B-scan with their foveal pit identified by red arrows, with supporting text to describe the detection in each case.
Figure S4: Examples of foveal pit detection for fovea-centred, cross-sectional OCT B-scans, with text overlaid to describe the detection process in each case.
Rater repeatability
@@
As a measure of rater repeatability, author J.B. re-selected the fovea coordinate for all 516 SLO images two months after the initial manual detection, and compared the $(x,y)$-positioning using mean absolute error and intra-class correlation. Like the training pipeline for the fovea detection model, fovea segmentation masks were generated for each batch of coordinates and compared using the area under the receiver operating characteristic curve (AUC) and dice coefficient. Evaluation was performed after all images were resized to a common image resolution of 768 $\times$ 768 pixels for interpretable comparison.
@@
The repeatability results are listed below in supplementary Table S1. Results for the i-Test cohort represent inter-rater repeatability (manual detection against accurate ground truth data from cross-referenced OCT B-scan using Choroidalyzer 16), and results for the remaining four cohorts of data represent intra-rater repeatability. Intra-rater repeatability had an excellent average intra-class correlation of 0.99 and a strong intra-class correlation for inter-rater repeatability of 0.82 was also reported.
@@
Others 0.96 0.93 0.99 0.99 3.39 3.00
i-Test 0.94 0.89 0.77 0.87 4.75 4.52
Table S1: Repeatability of detecting the fovea on the i-Test dataset (which measure inter-rater repeatability) and the remaining cohorts (which measure intra-rater repeatability). AUC, area under the receiver operating characteristic curve; ICC, intra-class correlation for single fixed raters; MAE, mean absolute error (in pixels)
@@
    Alternating rule: It’s common to see arteries and veins appear on the fundus in an alternative nature, for efficient oxygen transport 47.
@@
    Thickness rule: Veins are typically thicker than arteries 41.
@@
    Brightness rule: Veins are typically darker than arteries 48.
@@
Figure S5: An exemplary SLO image with annotations to describe each of the four rules for artery-vein classification.
A protocol was created to ensure consistent and comparable segmentations by each grader using ITK-Snap 23. After application of the binary vessel segmentation model to detect all the vessels in each SLO image, the graders followed the rules listed below during pixel-level annotation:
@@
    Remove false positive pixels, i.e. retinal tissue which has been classed as vessel.
@@
Figure S6: An exemplary SLO image (A) with a raw binary vessel mask (B) superimposed in green with annotations to describe each segmentation rule followed during annotation.
@@
Grader 1 vs. Grader 2 0.99 0.96 0.97 0.94
Grader 1 vs. Grader 3 0.99 0.95 0.96 0.95
Grader 2 vs. Grader 3 0.99 0.95 0.96 0.94
@@
Supplementary Table S3 show the qualitative grading results from the clinical ophthalmologist (I.M.) for the manual annotations. All SLO images were graded as having ‘good quality’ or above, except for one, and there was no annotation which received a bad or worse rating. Across the image quality categories, a total of 19 SLO images were marked as ‘very good’, 10 as ‘good’ and 1 as ‘okay’.
@@
One of our image graders (author J.B.) identified any major errors in the original segmentation labels and corrected them using ITK-Snap 23. Supplementary Fig. S8 shows an SLO image with the most significant error (A), the SLO image with the original labelling overlaid (B), and the SLO image with the corrected version overlaid (C).
@@
Figure S8: SLO image (A), original manual annotation superimposed 11 (B) and corrected manual annotation (C). Here, many of the vessels in the macula were all assumed to be venules in the original annotation.
Image Quality Original 11 Corrected Both Very good ($N=6$) 0 3 3 Good ($N=4$) 0 2 2 Okay ($N=10$) 0 7 3 Bad ($N=3$) 0 2 1 Total ($N=23$) 0 14 9 Label Artery-Vein classification Original 11 VG: 6, G: 8, O: 6, B: 2, VB: 1 Corrected VG: 10, G: 11, O: 2, B: 0, VB: 0
Table S4: Qualitative adjudication between the original and corrected manual annotations for artery and vein classification in the RAVIR dataset 11, from a clinical ophthalmologist (I.M.). VG, very good; G, good; O, okay; B, bad; VB, very bad.
@@
SLOctolyzer’s segmentation model architectures
Binary vessel detection
For the binary, all-vessel segmentation model, we use a custom UNet deep learning architecture with a depth of 4, as shown in supplementary Fig. S9. A convolution block consists of two $3\times 3$ convolution layers, each followed by batch normalisation 49 and non-linear rectified linear unit activation (ReLU). After an initial convolution block, the model’s DownBlocks double the channel dimension from 48 to 768 while decreasing the spatial dimension by $\nicefrac{{1}}{{2}}$ to a final, encoded block with a spatial resolution of 20 $\times$ 15 with 768 features maps. The model’s UpBlocks reverse the DownBlocks, with each block reducing the channel dimension by $\nicefrac{{1}}{{2}}$ after applying a $2\times 2$ transposed convolution layer with a $2\times 2$ stride which scales the spatial dimension by 2. A final $1\times 1$ output convolution layer reduces the channel dimension to 1, resulting in the probabilistic segmentation map as output
As described in the main text, offline augmentation extracts 20 random patches of size 320 $\times$ 240 pixels from each SLO image (of native image resolution 768 $\times$ 768 pixels) and these are used for training, as shown in the input/output stage of supplementary Fig. S9. At inference, the entire SLO image of image resolution 768 $\times$ 768 pixels is fed into the network, outputting a 768 $\times$ 768 pixel probability segmentation map.
Figure S9: Custom UNet model architecture for the binary, all-vessel segmentation model.
Fovea detection
The model used for fovea detection was an ‘off-the-shelf’, pre-trained model from SegmentationModelsPytorch 25 which utilised a MobileNetV3 29 backbone encoder arm with a UNet decoder arm (the default setup from SegmentationModelsPytorch). The exact Python code to instantiate the model is seen below, with the SegmentationModelsPytorch package assumed to be installed.
Listing 1: Python code to reproduce model for fovea detection.
import segmentation_models_pytorch as smp
model = smp.Unet(encoder_name="timm-mobilenetv3_small_075",
encoder_weights="imagenet",
in_channels=1,
classes=1)
Artery-vein-optic disc detection
The model used for artery-vein-optic disc detection was an ‘off-the-shelf’, pre-trained model from SegmentationModelsPytorch 25 which utilised a ResNet101 32 backbone encoder arm with a UNet decoder arm (the default setup from SegmentationModelsPytorch). The exact Python code to instantiate the model is seen below, with the SegmentationModelsPytorch package assumed to be installed.
Listing 2: Python code to reproduce model for artery-vein-optic disc detection.
import segmentation_models_pytorch as smp
model = smp.Unet(encoder_name="resnet101",
encoder_weights="imagenet",
in_channels=1,
classes=4)
@@
Supplementary Fig. S10 shows the key differences in global and local vessel calibre. Global vessel calibre is the ratio between vessel pixels and skeletonised vessel pixels (number of yellow pixels divided by brown pixels). This provides a coarse measure of calibre. Local vessel calibre measures the diameter of each vessel segment by traversing it’s length and measuring the diameter of the largest circle which fits entirely within it. A vessel segments’ calibre is then the average value of the list of circle diameters collected as the vessel segment is traversed. Local vessel calibre is thus the overall average of all vessel segment calibres, hence a more granular approach to assess vessel calibre.
Figure S10: (A) SLO binary vessel map. Binary vessel map with skeleton overlaid in brown (B) and individual vessel segments overlaid in an array of colours (C). Note that the optic disc area is removed from the vessel map during feature measurement.
@@
Supplementary Fig. S11 shows the core steps to run the SLOctolyzer pipeline. SLOctolyzer can be run either from an integrated development environment (IDE), such as Jupyter, Spyder or Visual Studio Code, or directly from the terminal. All that is required is the input/output directories in a simple configuration text file. A process log for the user is outputted as the pipeline processes each image in turn, utilising GPU acceleration if available. When inputting image files, an optional spreadsheet with additional metadata such as the eye laterality, location and transversal spatial sampling length-scale (in microns-per-pixel) can be provided, so that measurements in pixel units can be converted into microns. For .vol this is not necessary as SLOctolyzer can access the file metadata.
Figure S11: Summary of the process of running SLOctolyzer on a batch ($n=3$) of images. (A) Directory of input SLO images to be analysed, with an optional spreadsheet which allows manual data entry of transversal spatial sampling length-scale, laterality and location. (B) Spreadsheet with optional metadata to accompany SLO images during pipeline. Note that missing information is supported. (C) Configuration file specifying the file paths to images and where results will be saved. (D) Demonstration of running SLOctolyzer via JupyterLab or via the terminal.
@@
Figure S12: Summary of the output of running SLOctolyzer on a batch ($n=3$) of images. (A) Directory of output folders for each of the three image files analysed. (B) Output files for the ‘opticdisc_768’ exemplar SLO image, including the original SLO image, segmentation masks, composite image with segmentations superimposed and a results file with metadata and feature measurements. (C) A directory with composite images of SLO and segmentation masks per image file for quick inspection of segmentation quality. (D) A composite results file with the metadata and features of each image file processed, stored row-wise to facilitate downstream statistical analysis.
@@
Supplementary Fig. S13 shows two examples which had the second largest outliers for tortuosity density and large-vessel calibre (AVR/CRAE/CRVE). In supplementary Fig. S13(A), while there is no qualitative difference in the binary vessel segmentation, the error in tortuosity density is quite large. This is likely due to vessel disconnectedness between the segmentations (purple arrows). Nevertheless, the local vessel calibre and fractal dimension residuals are very low.